OpenStudy (cggurumanjunath):

integrate sin(x) / (5+3sinx)

OpenStudy (adamaero):

Are you in a certain chapter?

OpenStudy (shadowlegendx):

Integrate the sin of my tan booty cos it be on FIYAH

OpenStudy (jhannybean):

\[\int \frac{\sin(x)}{5+3\sin(x)}dx\] kind of resembles a complicated inverse tan function just by how it looks, that's a start.

OpenStudy (adamaero):

trig sub is not worken' out

OpenStudy (jhannybean):

Yeah, it might, something to do with setting \( u= \sin(x)\) possibly?

OpenStudy (jhannybean):

And then it'd look more like \[\int \frac{1}{1+x}dx = \tan^{-1}(x)\] or something like that. Haven't checked wolfram yet, tried analyzing the structure of the function first. Hmm

OpenStudy (adamaero):

CGGURUMANJUNATH can you give any more information, such as where is this from??

OpenStudy (cggurumanjunath):

quiz

OpenStudy (adamaero):

calc II or III?

OpenStudy (cggurumanjunath):